A NaOH solution has a density of 1.22 g/mL and is 20.0 wt % NaOH. What is the molality of the solution?
Answer #1

20wt%NaOH means 20g of NaOH present in 100g of solution

mass of solute( NaOH) = 20g

mass of solution = 100g

mass of solvent = mass of solution - mass of solute

                           = 100-20   = 80g

molar mass of NaOH = 40g/mole

molality   = mass of solute * 1000/gram molar mass of NaOH * mass of solvent in g

                = 20*1000/(40*80)

                = 6.25m   >>>>>answer